# Probing Language Models on Their Knowledge Source

### v1

@@ 3.3 Models
In our probing set-up (Section 4), we use the following activations: $\sigma(X^{(l)}_{mlp}W_{\textit{mlp}}^{(l)})$ the first layer of the MLP (referred as MLP-L1 in this paper), $\sigma(X^{(l)}_{mlp}W_{\textit{mlp}}^{(l)})W_{\textit{proj}}^{(l)}$ the output of the MLP (i.e. second layer, referred as MLP in this paper), and $A^{(l)}$ the output of the MHSA. We consider the first and second MLP layers activations, based on Geva et al. (2021) work, and also the MHSA activations as the attentions play a crucial role in information selection from the MLP memory Geva et al. (2023).
@@
Table 2: Activation dimensions for Phi-1.5, Pythia-1.4B, Llama3-8B and Mistral-7B for the different considered modules (MLP, MLP-L1 and MHSA)
@@ 5 Results and Discussion
In Figure 5, we can first observe that the features linked to $\bar{o}$, the subject $s_{q}$ and the relation $r_{q}$ exhibit a correlation with the used knowledge source for MLP and MLP-L1 activations. The most predictive features are those of $r_{q}$, i.e. the relation token in the query. Starting from the mid-layers of the relation token, the success rate increases significantly, reaching about 80%. This finding is consistent with prior research, which indicates that LLMs primarily store knowledge in the MLPs Meng et al. (2023); Geva et al. (2021). Moreover, it supports Geva et al. (2023)’s insights on the information extraction process, where the relation token retrieves attributes from $s_{q}$ (a process referred to as Attribute Extraction).
Additionally, it is noteworthy that the knowledge source can be detected directly starting from the perturbing object $\bar{o}$. This shows that detecting a potentially harmful conflict knowledge statement is possible early in the LLM inference process. MHSA activations are less connected to the used knowledge source than MLP and MLP-L1 activations.
The results of the control experiments conducted on the first token of the input indicate that the learned patterns in the object, subject, and relation are not arbitrary. The success rates of most LLMs for the first token appear to be random (about 0.5), with the exception of Pythia-1.4B, where the first token provides a slight indication of the knowledge source, although no significant fluctuations are observed.
@@
Figure 5: Performance of the linear classifier in identifying knowledge sources across different layers and modules (MLP, MLP-L1, MHSA). The plots show success rates for classifiers trained on activations from object, subject, and relation tokens, with the first token used as a control (see Section 4.1 for more details on the control experiment.) Results are reported for the Mistral-7B, Phi-1.5, Llama3-8B, and Pythia-1.4B models. Solid lines represent the average success rates across relation groups, while shaded areas denote the weighted standard error with a 95% confidence interval. See Section 4.3 for further details on the evaluation methodology.
@@ 8 Conclusion
In this study, we introduced a novel probing framework to investigate if we can detect when LLMs switch from PK to CK. Our findings reveal that specific model activations are significantly correlated with the used knowledge source. This opens the door for future work investigating the mechanism at play when such a switch occurs, and finally to building models that can better control this behavior.
@@ 11 Acknowledgements
We would like to thank BNP Paribas and the French National Association for Research and Technology (ANRT) who founded this project under the CIFRE program (2023/1673).
@@ Appendix A Data Characteristics
Appendix B Hardware and Software

### v3

@@ 3.3 Models
In our probing set-up (Section 4), we use the following activations: $\sigma(X^{(l)}_{mlp}W_{\textit{mlp}}^{(l)})$ the first layer of the MLP (referred as MLP-L1 in this paper), $\sigma(X^{(l)}_{mlp}W_{\textit{mlp}}^{(l)})W_{\textit{proj}}^{(l)}$ the output of the MLP (i.e. second layer, referred as MLP-L2 in this paper), and $A^{(l)}$ the output of the MHSA. We consider the first and second MLP layers activations, based on Geva et al. (2021) work, and also the MHSA activations as the attentions play a crucial role in information selection from the MLP memory Geva et al. (2023).
@@
Table 2: Activation dimensions for Phi-1.5, Pythia-1.4B, Llama3-8B and Mistral-7B for the different considered modules (MLP-L2, MLP-L1 and MHSA)
@@ 5 Results and Discussion
In Figure 5, we can first observe that the features linked to $\bar{o}$, the subject $s_{q}$ and the relation $r_{q}$ exhibit a correlation with the used knowledge source for MLP-L2, MLP-L1, and MHSA activations.
Our framework successfully transfer the learned knowledge source patterns from one relation group to another, generalizing well from one group to another. The most predictive features are those of $r_{q}$, i.e. the relation token in the query. On certain layers, the success rate increases significantly, reaching 87% for Pythia-1.4b on the 15th layer at the relation token position.
This finding is consistent with prior research, which indicates that LLMs primarily store knowledge in the MLPs Meng et al. (2023); Geva et al. (2021). Moreover, it supports Geva et al. (2023)’s insights on the information extraction process, where the relation token retrieves attributes from $s_{q}$’s MLPs using the MHSA (a process referred to as Attribute Extraction).
Additionally, it is noteworthy that the knowledge source can be detected directly starting from the perturbing object $\bar{o}$. This shows that detecting a potentially harmful conflict knowledge statement is possible early in the LLM inference process. MHSA activations are less connected to the used knowledge source than MLP-L2 and MLP-L1 activations.
The results of the control experiments conducted on the first token of the input indicate that the learned patterns in the object, subject, and relation are not arbitrary. The success rates of all LLMs for the first token appear to be random (about 0.5).
@@
Figure 5: Performance of the linear classifier in identifying knowledge sources across different layers and modules (MLP-L2, MLP-L1, MHSA). The plots show success rates for classifiers trained on activations from object, subject, and relation tokens, with the first token used as a control (see Section 4.1 for more details on the control experiment.) Results are reported for the Mistral-7B, Phi-1.5, Llama3-8B, and Pythia-1.4B models. Solid lines represent the average success rates across relation groups, while shaded areas denote the weighted standard error with a 95% confidence interval. See Section 4.3 for further details on the evaluation methodology.
@@ 8 Conclusion
In this study, we introduced a novel probing framework to investigate if we can detect when LLMs switch from PK to CK. Our findings reveal that specific model activations are significantly correlated with the used knowledge source with a success rate reaching 87% for Pythia-1.4b. Additionally, our framework is able to transfer the learned knowledge source patterns from one relation group to another. This opens the door for future work investigating the mechanism at play when using CK or PK, and finally to building models that can better control this behavior.
@@ 11 Acknowledgements
We would like to thank BNP Paribas and the French National Association for Research and Technology (ANRT) who founded this project under the CIFRE program (2023/1673). We also thank Etienne Boisseau for his help on this paper.
@@ Appendix A Data Characteristics
Appendix B Data Undersampling Seed Impact
To examine the impact of random seed selection in undersampling for balanced CK and PK classes, we conducted an experiment with various seeds to determine if seed choice influenced model performance. Changing the seed for the undersampling of the majority class introduces significant variations in the success rate of our classifiers. This effect can be explained by the fact that the minority class (PK) has fewer samples than the majority class (CK), meaning there are very few CK examples in common between the datasets generated by two different seeds. In some cases, we observed a standard deviation of up to 11 points of accuracy in the success rate when varying the seed. However, the results stayed consistent with our conclusions across all choices of seeds.
Appendix C Hardware and Software